# Upper critical field and pairing symmetry of Ising superconductors

### v1

@@ Abstract
The result to highlight here is that the $\sqrt{\lambda_{0}/\Delta_{0}}$-scaling only appears if the superconducting order parameter has a finite singlet gap at the SOC nodes. For a triplet order a $\lambda_{0}/\Delta_{0}$-scaling is regained (see [15] for the critical field obtained for different pairing symmetries on a $\Gamma$-pocket). The fact that the experimental data for monolayer NbSe2 and TaS2 are related by $\sqrt{\lambda_{0}/\Delta_{0}}$-scaling (Fig. 2) is therefore a strong indicator that the superconducting order is mainly a singlet at the SOC nodes.
Figure 3: Mono- and bi-layer NbSe2[6] and TaS2[8, 7]. For bilayers a fit for the interlayer bias potential $\delta\mu$ is shown. For multilayers a lower bound of $H_{c}(T)$ is Eq. (12), the suppression from hopping between the layers at $\delta\mu=0$. The upper bound considers the layers as uncoupled monolayers at the given $T_{c}$ ($\delta\mu\rightarrow\infty$).
Multilayer. In multilayer TMDs, individual layers break inversion symmetry, yet in the 2H-stacking the overall inversion is preserved, and the sign of the SOC alternates between layers. Layers are coupled by an interlayer hopping $t_{\perp}$. This coupling is stronger for the $\Gamma$-pocket[34] due to the increased presence of $d_{z^{2}}$-orbitals[35, 36]. Motivated by this, all pockets should be included also in the multilayer model.
For intralayer singlet superconductivity the susceptibility can never reach beyond the value of the normal state van Vleck-susceptibility[37, 9, 38], as the remaining Pauli susceptibility originates from the Fermi surface. In a bilayer 2H-TMD the SOC nodes disappear and instead two sets of degenerate bands are separated by a distance proportional to $t_{\perp}$. In our model we take the same $t_{\perp}$ for all pockets. The now leading term $\delta\chi^{j}_{0}$ is determined by the ratio $t_{\perp}/\lambda_{0}^{j}$. For $K$- and $\Gamma$-pockets respectively, we find
$\delta\chi^{K,\text{2L}}(T)\approx\frac{t_{\perp}^{2}}{t_{\perp}^{2}+\left(\lambda_{0}^{K}\right)^{2}}(\chi_{P}-\chi_{\text{sg}}(T)),$ (12)
$\delta\chi^{\Gamma,\text{2L}}(T)\approx\frac{t_{\perp}}{\sqrt{t_{\perp}^{2}+\left(\lambda_{0}^{\Gamma}\right)^{2}}}(\chi_{P}-\chi_{\text{sg}}(T)).$ (13)
For any number of layers $N$ the leading term is $\delta\chi^{j}_{0}$, for up to $N=5$ layers[15].
Another non-negligible factor in the multilayer case is a small inversion symmetry breaking coming from a substrate. In the bilayer, we hence include an interlayer bias potential[39], $\delta\mu$. At the FS this splits the spin-polarized bands, and the SOC nodes reemerge. The inversion symmetry broken bilayer can therefore be modeled as two monolayer bands with an effective SOC $\tilde{\lambda}_{0}$. From a linearization of the bands close to the SOC nodes,
$\tilde{\lambda}^{j}_{0}=\frac{\delta\mu}{\sqrt{\delta\mu^{2}+t_{\perp}^{2}}}\lambda^{j}_{0}.$ (14)
If $\delta\mu$ is sufficiently large, the critical field thus scales as $H_{c}^{2L}/H_{p}^{2L}\propto\sqrt{\frac{\tilde{\lambda}^{\Gamma}_{0}}{\Delta_{0}^{\Gamma,\text{2L}}}}$. More generally, compared to the monolayer, the bilayer $H_{c}$ depends on an increased number of parameters: $\lambda^{j}_{0},\Delta_{0}^{j,\text{2L}},t_{\perp},\delta\mu$ for each pocket $j$. In results of Fig. 3 the value $\delta\mu$ has been fitted to experimental data, while $\lambda^{j}_{0},t_{\perp}$ are given by Ref. [8], and $\Delta_{0}^{j,\text{2L}}$ by $T_{c}$ for the fitted dataset. For both NbSe2 and TaS2 the find $\delta\mu<10$meV, which is reasonable [40, 41, 39, 42, 43].
To deal with multiple model parameters, we consider an upper and lower bound for the $N$-layer $H_{c}$. The lower bound is given by Eq. (12), where the sample is considered fully isolated from its environment. The upper bound is the limit where each layer is treated as an isolated monolayer. For layers $N>2$, the bounds can be calculated (see [15] for details), while additional parameters are required to model the effects of a substrate.
In contrast to the monolayer case the multilayer $H_{c}$ does not carry a strong indication about the pairing symmetry at specific points in momentum space. However, for a fully triplet order the interlayer hopping would not suppress $H_{c}$[37, 9].
@@
Conclusions. We have demonstrated the necessity of including all bands when calculating the upper critical field of a superconductor. The critical field can be limited by the sections of the FS where $\bm{g}$ or $\bm{d}$ is no longer perpendicular to the field. We have illustrated this principle by comparing $H_{c}$ for Ising SOC superconductors either having or not pockets where the SOC has nodal points. We have shown a huge difference in $H_{c}$ arising for different pairing symmetries at these few SOC-node points. Therefore, the behavior of the upper critical field could be used to indicate which pairing symmetries at specific points in momentum space are favored.
@@
For monolayer 1H-NbSe2 and 1H-TaS2, our analysis shows that a dominant triplet order parameter is inconsistent with the experimental data. This rules out any purely $f$-wave[44, 45] or nematic order[46, 47, 48]. Furthermore, even though our calculation only includes a uniform singlet order, other proposed orders for NbSe2 are consistent with our results. This includes a nodal singlet, a topologically non-trivial singlet order such as a $d+id$-wave, or the recently proposed nodal topological superconductor[29, 44, 49, 50, 51]. A mixed parity order, $s+f$-wave[52, 53] or a $d_{y}$-triplet component appearing under a field[54, 10, 24], would also only affect the subleading order of $\delta\chi(T)$ and thus only slightly change the value of $H_{c}$. It should also be reiterated that the signature from the $K$-pockets are similarly almost entirely hidden, as their potential influence on $H_{c}$ is of the same size as disorder effects[11, 10].
@@
Other works have also found the $\Gamma$-pocket essential to accurately describe a pairing mechanism dependent on spin-fluctuations[48], reinforcing the conclusion that the nodal points can play a vital role.
@@
Note that a Rashba SOC $\lambda_{R}$, e.g., due to a substrate, lifts the Ising SOC nodes. We show that the correction to the susceptibility difference is only of order $(\lambda_{R}/\lambda_{0})^{2}$, and hence the influence on $H_{c}$ is negligible (see [15]).
@@
In this work, we focused on Ising SOC superconductors as characteristic inversion-symmetry-breaking materials with only a few points in the FS without a SOC splitting. However, the idea is general as it relies on symmetries and shows that the critical field could similarly be used to extract information on the nature of the order parameter in hetero- and/or Moiré structures.

### v2

@@ Abstract
The result to highlight here is that the $\sqrt{\lambda_{0}/\Delta_{0}}$-scaling only appears if the superconducting order parameter has a finite singlet gap at the SOC nodes. For a triplet order a $\lambda_{0}/\Delta_{0}$-scaling is regained (see [SupMat] for the critical field obtained for different pairing symmetries on a $\Gamma$-pocket). The fact that the experimental data for monolayer NbSe2 and TaS2 are related by $\sqrt{\lambda_{0}/\Delta_{0}}$-scaling (Fig. 2) is therefore a strong indicator that the superconducting order is mainly a singlet at the SOC nodes.
Figure 3: Prediction for the upper critical field in 1H-NbSe2[Rahn2012, SupMat] with a varying chemical potential $\mu$, corresponding to hole and electron doping in Eq. (12). $H_{p}$ is scaled with $T_{c}(\mu)$.
Experimental prediction. An experimental consequence of the scaling with SOC can be predicted for a doped monolayer. We note that when doping bilayer 2H-NbSe2 via gating, Ref. [Xi2016a] observed a large 30% change of $T_{c}$, hence gating should be a powerful probe. In our tight-binding modeling of monolayer NbSe2[Rahn2012], several parameters are affected by doping. Importantly, due to the symmetry conditions on the SOC, the effective SOC at the $\Gamma$-pocket decreases under electron doping while it increases around $K$: $\lambda_{0}^{j}(\mu)\approx\lambda_{0}^{j}(0)\left(1+\mu\alpha_{\lambda}^{j}\right)$, with $\alpha_{\lambda}^{\Gamma}>0$ and $\alpha_{\lambda}^{K}<0$ for electron doping $\mu<0$. Secondly, using the McMillan formula to relate $T_{c}$ and $\Delta_{0}$ as in the experimental analysis of Ref. [Xi2016a] (see [SupMat] for all details), we approximate the change of pairing at low doping, $\Delta_{0}(\mu)\approx\Delta_{0}(0)\left(1+\mu\alpha_{T}\right)$, with $|\alpha_{T}|<|\alpha_{\lambda}^{\Gamma}|,|\alpha_{\lambda}^{K}|$. Finally, we extract the evolution of the density of states for each pocket $N^{j}(\mu)$, as well as the $\lambda_{0}^{j}(\mu)$, from a detailed tight-binding model, but our main prediction remains for parabolic bands with the same shifts due to $\mu$ (see [SupMat]). Namely, for a singlet pairing on a FS with three pockets we find a doping-dependent critical field:
$\frac{H_{c2}(\mu)}{H_{p}}=\sqrt{\frac{N^{\Gamma}(\mu)+2N^{K}(\mu)}{N^{\Gamma}(\mu)\frac{\delta\chi^{\Gamma}(T,\mu)}{\chi_{P}-\chi_{\text{sg}}(T)}+2N^{K}(\mu)\frac{\delta\chi^{K}(T,\mu)}{\chi_{P}-\chi_{\text{sg}}(T)}}},$ (12)
that crucially has the slope $\partial\frac{H_{c2}}{H_{p}}/\partial\mu\propto(\alpha_{\lambda}^{j}-\alpha_{T})$ having opposite sign for when the $\Gamma$-pocket is included or not, as seen in Fig. 3. Hence, also for a realistic bandstructure, we predict a linearly decreasing $H_{c2}/H_{p}$ with small electron doping [^7] Electron doping of NbSe2 is possible with intercalation, field effect transistors, and misfit structures. Hole doping, while possible, is more difficult to achieve for large doping values, due to the high work function in TMDs[Zullo2023]., or correspondingly increasing with hole doping, as a clear signature of the $\sqrt{\lambda_{0}^{\Gamma}/\Delta_{0}}$-scaling due to the dominant role of the $\Gamma$-pocket, while a linearly increasing trend is predicted in absence of the $\Gamma$-pocket.
@@
Discussion and conclusions. We have demonstrated the necessity of including all bands when calculating the upper critical field of a superconductor. The critical field can be limited by the sections of the FS where $\bm{g}$ or $\bm{d}$ is no longer perpendicular to the field. We have illustrated this principle by comparing $H_{c2}$ for Ising SOC superconductors either having or not pockets where the SOC has nodal points. We have shown a huge difference in $H_{c2}$ arising for different pairing symmetries at these few SOC-node points. Therefore, the behavior of the upper critical field could be used to indicate which pairing symmetries at specific points in momentum space are favored.
@@
For monolayer 1H-NbSe2 and 1H-TaS2, our analysis shows that a dominant triplet order parameter is inconsistent with the experimental data. This rules out any purely $f$-wave[Shaffer2020] or nematic order[Cho2022, Roy2025, Siegl2024]. Furthermore, even though our calculation only includes a uniform singlet order, other proposed orders for NbSe2 are consistent with our results. This includes a nodal singlet, a topologically non-trivial singlet order such as a $d+id$-wave, or the recently proposed nodal topological superconductor[He2018, Shaffer2020, Cohen2024, Margalit2021]. A mixed parity order, $s+f$-wave[Das2023, Wan2022] or a $d_{y}$-triplet component appearing under a field[Tang2021, Mockli2020, Kuzmanovi2022], would also only affect the subleading order of $\delta\chi(T)$ and thus only slightly change the value of $H_{c2}$. It should also be reiterated that the signature from the $K$-pockets are similarly almost entirely hidden, as their potential influence on $H_{c2}$ is of the same size as disorder effects[Ilic2017, Mockli2020].
@@
Other works have also found the $\Gamma$-pocket essential to accurately describe a pairing mechanism dependent on spin-fluctuations[Siegl2024], reinforcing the conclusion that the nodal points can play a vital role.
Further, our conclusions are consistent with the available data on few-layer TMDs with 2H-stacking [^8] This will be discussed in detail elsewhere.. In homobilayers, the overall inversion symmetry is preserved. However, the SOC within each layer has nodal points which should be taken into account when calculating $H_{c2}$.
@@
Note that a Rashba SOC $\lambda_{R}$, e.g., due to a substrate, lifts the Ising SOC nodes. We show that the correction to the susceptibility difference is only of order $(\lambda_{R}/\lambda_{0})^{2}$, and hence the influence on $H_{c2}$ is negligible (see [SupMat]).
@@
In this work, we focused on monolayer Ising SOC superconductors as characteristic inversion-symmetry-breaking materials with only a few points in the FS without a SOC splitting. However, the idea is general as it relies on symmetries and shows that the critical field could similarly be used to extract information on the nature of the order parameter in hetero- and/or Moiré structures.
Finally, our approach might be used to investigate the nature of pairing in Ising superconductors whose bandstructure and/or superconductivity has been altered by a wide range of experimental control parameters, such as pressure or strain ([Abdel-Hafiez2016, Henriquez-Guerra2024]), electron- and hole-doping via field-effect gating or the introduction of an external electric field[Liu2017, Ye2012]. Notably, the family of misfit dichalcogenide compounds, formed by the stacking of rocksalt and TMD layers, offers an efficient way to tune the effective TMD chemical potential, potentially to large values.